# A Unified Approach for Multi-granularity Search over Spatial Datasets

### v2

@@ VII-C Efficiency of Dataset Search
VII-C1 Top-$k$ Overlap-based Query
@@
VII-C2 Top-$k$ Point-wise Query
@@
Effect of Top-$k$ Search.  1) Exact Hausdorff. We first show the top-$k$ Hausdorff search performance of our ExactHaus against the ScanHaus [12] and the IncHaus [46]. The results are shown in Figure 14. With the increase of $k$, we can observe that our ExactHaus based on the unified pruning mechanism drastically improves the performance of top-$k$ Hausdorff search compared with the ScanHaus and IncHaus, achieving up to at most three orders of magnitude speed improvements against the ScanHaus and IncHaus. This proves that our fast bound estimation technology is effective.
2) Approximate Hausdorff. The approximate Hausdorff is further evaluated with decreasing error threshold $\epsilon$. Since the error threshold increases as the resolution decreases, we present the experimental results of our proposed ApproHaus in pairwise Hausdorff computation (i.e., computing the Hausdorff distance between two spatial datasets) and top-$k$ Hausdorff search as the resolution decreases in Figure 15. We observe that the runtime of pairwise ApproHaus computation degrades slightly with the increase of the error threshold. This is because in order to achieve a smaller error, the algorithm needs to scan more child nodes in the dataset index for approximate Hausdorff computation and terminate later.
@@
Effect of Outlier Removal.  Figure 17 shows the experimental results of top-$k$ Hausdorff search under three conditions: without using outlier removal (Origin), using our ExactHaus, and utilizing the SOTA INNE [13, 80] for outlier removal. The results from INNE are set as the ground truth. Also, we compare the execution time of these two outlier removal algorithms. The results demonstrate that our outlier removal technique can effectively remove outliers in the dataset by detecting leaf nodes with excessively large radii, thereby enhancing the accuracy of the search (up to 90%). Moreover, our algorithm exhibits significantly faster performance, surpassing INNE by five orders of magnitude, due to the substantial reduction in data points achieved by detecting kneedle points.
Effect of Node Capacity $f$.  Figure 19 shows the runtime of the three index-based methods when computing the Hausdorff distance between a pair of datasets. We observe that as the capacity of leaf nodes increases, the execution time of the pairwise Hausdorff gradually increases. This is because larger leaf nodes imply that more data points need to be scanned one by one, which affects the performance. Figure 19 shows that as the capacity of leaf nodes $f$ increases, the running time of the three methods does not change much as the capacity of leaf node $f$ increases, which indicates that the impact of node capacity on the Top-$k$ queries is minimal. This is because more time is spent on pruning, sorting, and computing bounds, rather than scanning the data points within the nodes.
@@
VII-D Efficiency of Data Point Search
@@
Nearest Neighbor Point Search.  We mainly compare our NNP search operation presented in Section VI-B against the k-nearest neighbor (kNN[59] that works similarly to the Hausdorff distance [46]. Figure 23 shows that our method can accelerate the NNP operation significantly, which achieves up to seven orders of magnitude speedup over kNN. This is because the kNN shares a common procedure with NNP to find the nearest points for points in the query datasets, and we can save much time by continuously traversing the unified index.
VII-E Case Study
Fig. 23 shows the visualization of the results that conduct trajectory search and analysis based on the Spadas system developed in this paper. Specifically, we added a transportation dataset collection to Spadas that includes a variety of trajectory datasets from Washington, D.C. An analyst aims to find $k$ trajectory datasets that are closely similar to a given query dataset from the Spadas to perform trajectory analysis tasks, such as trajectory near-duplicate detection [69] and traffic congestion prediction [81, 83].

### v3

@@ 7.3. Efficiency of Dataset Search
7.3.1. Top-$k$ Overlap-based Query
@@
7.3.2. Top-$k$ Point-wise Query
@@
Effect of Top-$k$ Search.  1) Exact Hausdorff. We compare the top-$k$ Hausdorff search performance of our ExactHaus with ScanHaus and IncHaus as $k$ increases. As shown in Figure 14, we can observe that our ExactHaus based on the unified pruning mechanism drastically improves the performance of top-$k$ Hausdorff search, achieving up to at most three orders of magnitude speed improvements against the ScanHaus and IncHaus. This proves that our fast bound estimation technology is effective.
2) Approximate Hausdorff. The approximate Hausdorff is further evaluated by decreasing the error threshold $\epsilon$. Since the error threshold increases as the resolution decreases, we present the experimental results of our proposed ApproHaus in pairwise Hausdorff computation (i.e., computing the Hausdorff distance between two spatial datasets) and top-$k$ Hausdorff search, as the resolution decreases, in Figure 15. We observe that the runtime of pairwise ApproHaus computation degrades slightly with the increase of the error threshold. This is because in order to achieve a smaller error, the algorithm needs to scan more child nodes in the dataset index for approximate Hausdorff computation and terminate later.
@@
(a) Accuracy of Top-$k$ Hausdorff Search
(b) Runtime of Top-$k$Hausdorff Search
Figure 16. Accuracy and runtime of three exemplar searches.
(a) Accuracy of Outlier Removal
(b) Runtime of Outlier Removal
Figure 17. Accuracy and runtime of top-$k$ Hausdorff search.
Figure 18. Capacity’s effects on pairwise Hausdorff search.
Figure 19. Index capacity’s effects on top-$k$ Hausdorff search.
(a) Runtime of overlap-based queries
(b) Runtime of point-wise query
Figure 20. Spatial dimension $d$’s effect on the Chicago dataset.
Figure 21. Range-based data point search as $R$ increases.
Figure 22. Nearest neighbor point search as $s$ increases.
Figure 23. Case study on trajectory datasets.
Effect of Outlier Removal.  Figure 17 shows the experimental results of top-$k$ Hausdorff search under three conditions: without using outlier removal (Origin), using our ExactHaus, and utilizing the SOTA INNE (bandaragoda2018isolation, ; zhao2019pyod, ) for outlier removal. The results from INNE are set as the ground truth. Also, we compare the execution time of these two outlier removal algorithms. The results demonstrate that our outlier removal technique can effectively remove outliers in the dataset by detecting leaf nodes with excessively large radii, thereby enhancing the accuracy of the search (up to 90%). Moreover, our algorithm exhibits significantly faster performance, surpassing INNE by five orders of magnitude, due to the substantial reduction in data points achieved by detecting kneedle points.
Effect of Node Capacity $f$.  Figure 19 shows the runtime of the three index-based methods when computing the Hausdorff distance between a pair of datasets. We observe that as the $f$ increases, the execution time of the pairwise Hausdorff gradually increases. This is because larger leaf nodes imply that more data points need to be scanned one by one, which affects the performance. Figure 19 shows that as the capacity of leaf nodes $f$ increases, the running time of the three methods does not change much, indicating that the impact of node capacity on the top-$k$ queries is minimal. This is because more time is spent on pruning, sorting, and computing bounds, rather than scanning the data points within the nodes.
@@
7.4. Efficiency of Data Point Search
@@
Nearest Neighbor Point Search.  We mainly compare our NNP search operation presented in Section 6.2 against the k-nearest neighbor (kNN(Taha2015, ) that works similarly to the Hausdorff distance (Nutanong2011, ). Figure 23 shows that our method can accelerate the NNP operation significantly, which achieves up to seven orders of magnitude speedup over kNN. This is because the kNN shares a common procedure with NNP to find the nearest points for points in the query datasets, and we can save much time by continuously traversing the unified index.
7.5. Case Study
Fig. 23 shows the visualization of the results that conduct trajectory search and analysis based on the Spadas system developed in this paper. Specifically, we added a transportation dataset collection to Spadas that includes a variety of trajectory datasets from Washington, D.C. An analyst aims to find $k$ trajectory datasets that are closely similar to a given query dataset from the Spadas to perform trajectory analysis tasks, such as trajectory near-duplicate detection (XiaoWLYW11, ) and traffic congestion prediction (ZhengZYSZ14, ; ZhouTWN13, ).